OpenStudy (anonymous):

Please help me with algebra 2!!!

OpenStudy (anonymous):

\[\frac{ 3 }{ \frac{ 3-x }{ x } +1}\]

OpenStudy (anonymous):

Help me simplify!!!

OpenStudy (anonymous):

\[\frac{3}{\frac{3-x}{x}+1}=\frac{3}{\frac{3}{x}}=x \]

OpenStudy (misty1212):

\[\frac{ 3 }{ \frac{ 3-x }{ x } +1}=\frac{ 3 }{ \frac{ 3-x }{ x } +1}\times \frac{x}{x}=\frac{3x}{3-x+x}\] \[=\frac{3x}{3}=x\]